Problem 7. Two identical conducting small spheres are placed with their centers r = 0.300 m apart. One is given a charge

of q1 = 12.0 nC and the other a charge of q2 = −18.0 nC. (a) Find the electric force exerted by one sphere on the other.

(b) Next, the spheres are connected by a conducting wire. Find the electric force between the two after they have come to

equilibrium.


1
Expert's answer
2021-07-23T16:45:01-0400
"q_1=12nC,q_2=-18nC,r=0.3m"

Force


"F=\frac{kq_1q_2}{r^2}=-\frac{9\times10^9\times12\times10^{-9}\times18\times10^{-9}}{0.3^2}=-2.16\times10^{-05}N"

Attractive force

(B)

the spheres are connected by a conducting wire

"q_1+q_2=12nC-18nC=-6nC"

"q_1=q_2=-3nC"

"F=\frac{kq_1q_2}{r^2}"


"F=\frac{9\times10^9\times3\times10^{-9}\times3\times10^{-9}}{0.3^2}=9\times10^{-07}N"

Repulsive force